OpenStudy (anonymous):

What is 4 1/2 - 3 5/8

OpenStudy (anonymous):

Give me a second sweetheart

OpenStudy (anonymous):

Okay:)

OpenStudy (anonymous):

7/8. Do you want me to explain it to you? I'm in flvs also

OpenStudy (anonymous):

Sure. I am Russian and I just start to take algebra classes.

OpenStudy (anonymous):

okay. I'll show you the work

OpenStudy (anonymous):

Okay.

OpenStudy (wolf1728):

Change 4 1/2to eighths 4 1/2 = 36/8 3 5/8 = 29/8 36/8 - 29/8 = 7/8

OpenStudy (anonymous):

Thank you. What operation must be some to both sides of the equation to solve p - -3 7/9 = 5/6? Add 5/6 Subtract 3 7/9 Subtract 5/6 Add 3 7/9 @wolf1728

OpenStudy (wolf1728):

Add 3 7/9

OpenStudy (anonymous):

\[Solve -12 + x < (with a line under \it) -36 for x\]

OpenStudy (anonymous):

I don't know how to do that.

OpenStudy (wolf1728):



OpenStudy (anonymous):

What is the answer?

OpenStudy (wolf1728):

I don't think you can solve for x because it isn't an equation. There's no equals sign.

OpenStudy (anonymous):

-12 + x < (with a line under <) - 36 for x

OpenStudy (wolf1728):

I can't solve that. It's like asking solve for x 7x + 3 How can you solve that?

OpenStudy (anonymous):

Oh I got it. Thank you fr trying to help.:-)

OpenStudy (wolf1728):

okay and you're welcome

OpenStudy (anonymous):

Could you help with more?

OpenStudy (anonymous):

Which property would you use to solve the equation f/3 = 0.27

OpenStudy (anonymous):

@calculusxy could you help with the last question.

OpenStudy (calculusxy):

What question would that be, may I ask?

OpenStudy (anonymous):

Which property would you use to solve the equation f/3 = 0.27

OpenStudy (calculusxy):

You can use the multiplication property of equality (I think that's how it's said). Meaning you multiply both sides by 3.

OpenStudy (anonymous):

Thank you.

OpenStudy (calculusxy):

Welcome :)